OpenStudy (nessapetrenko):

Solve for x. 3(-2x - 1) = 2(x + 5) A) - 13 8 B) - 3 4 Eliminate C) 13 8 D) 3 4

OpenStudy (anonymous):

-6x-3=2x+10 -13=8x x=-13/8

OpenStudy (nessapetrenko):

Thanks @BatistaFilho